# Bell Nonlocality from Wigner Negativity in Qudit Stabilizer States

### v3

@@ Acknowledgements.
We acknowledge funding from Horizon Europe Research and Innovation Actions under Grant Agreement 101080173 CLUSTEC and from the Plan France 2030 through the ANR-22-PETQ-0006NISQ2LSQ and ANR-22-PETQ-0007 EPiQ projects.
@@ Appendix A Wigner Negativity
For any bipartite state $\sigma=\sum_{i,j,k,l}\sigma_{ijkl}|i\rangle\langle j|\otimes|k\rangle\langle l|$, the Bell operator’s expectation value from Eq. (16) is
$\mathrm{Tr}\left(\mathcal{B}_{\nu}\sigma\right)$ $=\sum_{u_{1},u_{2}\in\mathbb{Z}_{d}^{2}}W_{u_{1},u_{2}}(|\Phi_{\nu}\rangle\langle\Phi_{\nu}|)\sum_{\begin{subarray}{c}k,l,\\ m,n\in\mathbb{Z}_{d}\end{subarray}}\sigma_{klmn}\langle l|U^{\phantom{\dagger}}_{\nu}A_{u_{1}}U^{\dagger}_{\nu}|i\rangle\langle n|A_{u_{2}}|m\rangle$
$=\frac{1}{d^{5}}\sum_{\begin{subarray}{c}(u_{1})_{z},(u_{2})_{z},\\ u_{x},j\in\mathbb{Z}_{d}\\ v_{1},v_{2}\in\mathbb{Z}_{d}^{2}\end{subarray}}\omega^{((u_{1})_{z}+(u_{2})_{z})j+\nu_{u_{x}+2^{-1}j}-\nu_{u_{x}-2^{-1}j}+(u_{1})_{z}(v_{1})_{x}+(u_{2})_{z}(v_{2})_{x}-u_{x}((v_{1})_{z}+(v_{2})_{z})}$
$\hskip 66.86414pt\cdot\sum_{\begin{subarray}{c}k,l,\\ m,n\in\mathbb{Z}_{d}\end{subarray}}\sigma_{klmn}\,\omega^{2^{-1}(v_{1})_{x}(v_{1})_{z}+2^{-1}(v_{2})_{x}(v_{2})_{z}}\langle l|U^{\phantom{\dagger}}_{\nu}X^{(v_{1})_{x}}Z^{(v_{1})_{z}}U^{\dagger}_{\nu}|k\rangle\langle n|X^{(v_{2})_{x}}Z^{(v_{2})_{z}}|m\rangle$
$=\frac{1}{d^{3}}\sum_{\begin{subarray}{c}u_{x},j\in\mathbb{Z}_{d}\\ v_{1},v_{2}\in\mathbb{Z}_{d}^{2}\end{subarray}}\delta_{-j,(v_{1})_{x}}\delta_{-j,(v_{2})_{x}}\omega^{\nu_{u_{x}+2^{-1}j}-\nu_{u_{x}-2^{-1}j}-u_{x}((v_{1})_{z}+(v_{2})_{z})+2^{-1}(v_{1})_{x}(v_{1})_{z}+2^{-1}(v_{2})_{x}(v_{2})_{z}}$
$\hskip 71.13188pt\cdot\sum_{\begin{subarray}{c}k,l,\\ m,n\in\mathbb{Z}_{d}\end{subarray}}\sigma_{klmn}\omega^{k(v_{1})_{z}+m(v_{2})_{z}+\nu_{k+(v_{1})_{x}}-\nu_{k}}\langle l|k+(v_{1})_{x}\rangle\langle n|m+(v_{2})_{x}\rangle$
$=\frac{1}{d^{3}}\sum_{\begin{subarray}{c}u_{x},j\in\mathbb{Z}_{d}\\ (v_{1})_{z},(v_{2})_{z}\in\mathbb{Z}_{d}\end{subarray}}\omega^{\nu_{u_{x}+2^{-1}j}-\nu_{u_{x}-2^{-1}j}-u_{x}((v_{1})_{z}+(v_{2})_{z})-2^{-1}j(v_{1})_{z}-2^{-1}j(v_{2})_{z}}$
$\hskip 78.24507pt\cdot\sum_{\begin{subarray}{c}k,l,\\ m,n\in\mathbb{Z}_{d}\end{subarray}}\sigma_{klmn}\omega^{k(v_{1})_{z}+m(v_{2})_{z}+\nu_{k-j}-\nu_{k}}\langle l|k-j\rangle\langle n|m-j\rangle$
$=\frac{1}{d}\sum_{u_{x},j\in\mathbb{Z}_{d}}\sum_{\begin{subarray}{c}k,l,\\ m,n\in\mathbb{Z}_{d}\end{subarray}}\sigma_{klmn}\omega^{\nu_{u_{x}+2^{-1}j}-\nu_{u_{x}-2^{-1}j}+\nu_{k-j}-\nu_{k}}\delta_{u_{x}+2^{-1}j,k}\delta_{u_{x}+2^{-1}j,m}\delta_{l,k-j}\delta_{n,m-j}$
$=\frac{1}{d}\sum_{u_{x},j\in\mathbb{Z}_{d}}\sigma_{u_{x}+2^{-1}j,u_{x}-2^{-1}j,u_{x}+2^{-1}j,u_{x}-2^{-1}j}\omega^{\nu_{u_{x}+2^{-1}j}-\nu_{u_{x}-2^{-1}j}+\nu_{u_{x}-2^{-1}j}-\nu_{u_{x}+2^{-1}j}}$
$=\frac{1}{d}\sum_{n,m\in\mathbb{Z}_{d}}\sigma_{n,m,n,m}=\langle\Phi|\sigma|\Phi\rangle\,.$ (26)
@@ Appendix B Relation to the CHSH inequality
with the Pauli operators $\sigma_{(0,0)}=\mathds{1}$, $\sigma_{(1,0)}=X$, $\sigma_{(1,1)}=Y$, $\sigma_{(0,1)}=Z$. The evaluations of the expectation values and nonlocal violations for the qudit Bell operators in Eq. (16) and Eq. (20) do not straightforwardly generalize to qubits using the above Wigner function. The reason is that the assignments of an lhv model are different from qudit Pauli operators to qubit Pauli operators if one measures a complete set of operators generating mutually unbiased bases. However, in case where $U_{\nu}=T$, the equivalent Bell operator for qubits in Eq. (16) is
$\mathcal{B}^{\phantom{\prime}}_{2}=\mathds{1}\otimes\mathds{1}+\sigma^{z}\otimes\sigma^{z}-\frac{1}{\sqrt{2}}\Big{(}T\sigma^{y}T^{\dagger}\otimes\sigma^{y}+T\sigma^{x}T^{\dagger}\otimes\sigma^{y}+T\sigma^{y}T^{\dagger}\otimes\sigma^{x}-T\sigma^{x}T^{\dagger}\otimes\sigma^{x}\Big{)}\,,$ (29)
@@ Appendix C Bounds on Gross’ Wigner function
$\mathrm{Tr}\left(\sigma\mathcal{B}\right)$ $=d^{2}\sum_{u_{1},u_{2}\in\mathbb{Z}_{d}^{2}}W^{\nu}_{u_{1},u_{2}}W_{u_{1},u_{2}}\left((U_{\nu}\otimes\mathds{1})^{\dagger}\sigma\,(U_{\nu}\otimes\mathds{1})\right)$
$\leq d^{2}\max_{u_{1},u_{2}\in\mathbb{Z}_{d}^{2}}|W^{\nu}_{u_{1},u_{2}}|\sum_{u_{1},u_{2}\in\mathbb{Z}_{d}^{2}}|W_{u_{1},u_{2}}\left((U_{\nu}\otimes\mathds{1})^{\dagger}\sigma\,(U_{\nu}\otimes\mathds{1})\right)|\,.$ (32)
Appendix D Bell Inequality for Stabilizer States
For the Bell operator’s expectation value in Eq. (24) of an arbitrary quantum state on $n$ qudits, $\rho=\sum_{\mathbf{r}\in\mathbb{Z}_{d}^{2n}}W_{\mathbf{r}}(\rho)A_{\mathbf{r}}$, we determine
$\sum_{\mathbf{r}\in\mathbb{Z}_{d}^{2n}}W_{\mathbf{r}}(\rho)\mathrm{Tr}\left(\mathcal{B}_{S}A_{\mathbf{r}}\right)$ $=\sum_{\mathbf{r}\in\mathbb{Z}_{d}^{2n}}W_{\mathbf{r}}(\rho)\sum_{\begin{subarray}{c}{\bm{u}}\in\Sigma\\ z\in\mathbb{Z}_{d}\end{subarray}}W^{\nu}_{{\bm{u}}}\,\mathrm{Tr}\left(A_{\mathbf{r}}U_{1}\,A_{{\bm{u}}+(0,z)_{1}}\,(U_{1})^{\dagger}\right)$
$=d^{-2n}\sum_{\mathbf{r}\in\mathbb{Z}_{d}^{2n}}W_{\mathbf{r}}(\rho)\sum_{\begin{subarray}{c}\mathbf{s},{\bm{u}}\in\Sigma\\ z\in\mathbb{Z}_{d}\end{subarray}}\mathrm{Tr}\left(U_{1}A_{\mathbf{s}}(U_{1})^{\dagger}A_{{\bm{u}}}\right)\mathrm{Tr}\left(A_{\mathbf{r}}U_{1}\,A_{{\bm{u}}+(0,z)_{1}}\,(U_{1})^{\dagger}\right)\,,$
with $\mathrm{Tr}\left(A_{{\bm{u}}}A_{{\bm{v}}}\right)=\delta_{{\bm{u}},{\bm{v}}}$, and
$\mathrm{Tr}\left(UA_{s}U^{\dagger}A_{u}\right)$ $=\frac{1}{d^{2}}\sum_{t\in\mathbb{Z}_{d}^{2}}\omega^{[s,t]+2^{-1}t_{x}t_{z}}\sum_{k,l\in\mathbb{Z}_{d}}\omega^{\nu_{k+t_{x}}-\nu_{k}-kl}\mathrm{Tr}\left(X^{t_{x}}Z^{t_{z}+l}A_{u}\right)$
$=\frac{1}{d^{2}}\sum_{t\in\mathbb{Z}_{d}^{2}}\omega^{[s,t]+2^{-1}t_{x}t_{z}}\sum_{k,l\in\mathbb{Z}_{d}}\omega^{\nu_{k+t_{x}}-\nu_{k}-kl+[(t_{x},t_{z}+l),u]-2^{-1}t_{x}(t_{z}+l)}$
$=\frac{1}{d^{2}}\sum_{t\in\mathbb{Z}_{d}^{2}}\sum_{k,l\in\mathbb{Z}_{d}}\omega^{[s,t]+\nu_{k+t_{x}}-\nu_{k}-kl+[(t_{x},t_{z}+l),u]-2^{-1}t_{x}l}$
$=\frac{1}{d}\sum_{t\in\mathbb{Z}_{d}^{2}}\sum_{k\in\mathbb{Z}_{d}}\omega^{[s,t]+\nu_{k+t_{x}}-\nu_{k}+[(t_{x},t_{z}),u]}\delta_{k,u_{x}-2^{-1}t_{x}}$
$=\delta_{s_{x},u_{x}}\sum_{t_{x}\in\mathbb{Z}_{d}}\omega^{\nu_{u_{x}+2^{-1}t_{x}}-\nu_{u_{x}-2^{-1}t_{x}}+(s_{z}-u_{z})t_{x}}\,,$
using $\mathrm{Tr}\left(X^{x}Z^{z}A_{(u_{x},u_{z})}\right)=\omega^{zu_{x}-xu_{z}-2^{-1}xz}$. Then, it is
$\phantom{=}d^{-2n}\sum_{\mathbf{r}\in\mathbb{Z}_{d}^{2n}}W_{\mathbf{r}}(\rho)\sum_{\begin{subarray}{c}\mathbf{s},{\bm{u}}\in\Sigma\\ z\in\mathbb{Z}_{d}\end{subarray}}\mathrm{Tr}\left(UA_{\mathbf{s}}U^{\dagger}A_{{\bm{u}}}\right)\mathrm{Tr}\left(A_{\mathbf{r}}U\,A_{{\bm{u}}+(0,z)_{1}}\,U^{\dagger}\right)$
$=\sum_{\mathbf{r}\in\mathbb{Z}_{d}^{2n}}W_{\mathbf{r}}(\rho)\sum_{\begin{subarray}{c}\mathbf{s},{\bm{u}}\in\Sigma\\ z\in\mathbb{Z}_{d}\end{subarray}}\delta_{(s_{1})_{x},(u_{1})_{x}}\delta_{(u_{1})_{x},(r_{1})_{x}}\delta_{\mathbf{s}_{\setminus s_{1}},{\bm{u}}_{\setminus u_{1}}}\delta_{\mathbf{r}_{\setminus r_{1}},{\bm{u}}_{\setminus u_{1}}}$
$\hskip 71.13188pt\sum_{k,m\in\mathbb{Z}_{d}}\omega^{\nu_{(u_{1})_{x}+2^{-1}k}-\nu_{(u_{1})_{x}-2^{-1}k}+((s_{1})_{z}-(u_{1})_{z})k+\nu_{(r_{1})_{x}+2^{-1}m}-\nu_{(r_{1})_{x}-2^{-1}m}+((u_{1})_{z}+z-(r_{1})_{z})m}$
$=d^{-1}\sum_{\begin{subarray}{c}\mathbf{r}_{\setminus(r_{1})_{z}}\in\Sigma\\ (r_{1})_{z}\in\mathbb{Z}_{d}\end{subarray}}W_{\mathbf{r}}(\rho)\sum_{(s_{1})_{z},k,m,z\in\mathbb{Z}_{d}}\delta_{k,m}\,\omega^{\nu_{(r_{1})_{x}+2^{-1}k}-\nu_{(r_{1})_{x}-2^{-1}k}+(s_{1})_{z}k+\nu_{(r_{1})_{x}+2^{-1}m}-\nu_{(r_{1})_{x}-2^{-1}m}+(z-(r_{1})_{z})m}$
$=\sum_{\begin{subarray}{c}\mathbf{r}_{\setminus(r_{1})_{z}}\in\Sigma\\ (r_{1})_{z}\in\mathbb{Z}_{d}\end{subarray}}W_{\mathbf{r}}(\rho)\sum_{(s_{1})_{z},k\in\mathbb{Z}_{d}}\omega^{\nu_{(r_{1})_{x}+2^{-1}k}-\nu_{(r_{1})_{x}-2^{-1}k}+\nu_{(r_{1})_{x}+2^{-1}k}-\nu_{(r_{1})_{x}-2^{-1}k}+((r_{1})_{z}+(s_{1})_{z})k}\delta_{k}$
$=d\sum_{\begin{subarray}{c}\mathbf{r}_{\setminus(r_{1})_{z}}\in\Sigma\\ (r_{1})_{z}\in\mathbb{Z}_{d}\end{subarray}}W_{\mathbf{r}}(\rho)\,.$

### v4

@@ Acknowledgements.
We sincerely thank Mark Howard for valuable insights. We acknowledge funding from Horizon Europe Research and Innovation Actions under Grant Agreement 101080173 CLUSTEC and from the Plan France 2030 through the ANR-22-PETQ-0006NISQ2LSQ and ANR-22-PETQ-0007 EPiQ projects.
@@ Appendix A Wigner Negativity
For any bipartite state $\rho$, the Bell operator’s expectation value from Eq. (9) is
$\mathrm{Tr}\left(\mathcal{B}_{\nu}\rho\right)$ $=\sum_{u_{1},u_{2}\in\mathbb{Z}_{d}^{2}}W_{u_{1},u_{2}}(|\Phi_{\nu}\rangle\langle\Phi_{\nu}|)\mathrm{Tr}\left(U^{\phantom{\dagger}}_{\nu}A_{u_{1}}U^{\dagger}_{\nu}\otimes A_{u_{2}}\rho\right)$ (29)
$=\mathrm{Tr}\left(\left(U^{\phantom{\dagger}}_{\nu}\otimes\mathds{1}\right)|\Phi_{\nu}\rangle\langle\Phi_{\nu}|\left(U^{\dagger}_{\nu}\otimes\mathds{1}\right)\rho\right)$ (30)
$\langle\Phi|\rho|\Phi\rangle$ (31)
@@ Appendix B Relation to the CHSH inequality
with the Pauli operators $\sigma_{(0,0)}=\mathds{1}$, $\sigma_{(1,0)}=\sigma_{x}$, $\sigma_{(1,1)}=\sigma_{y}$, $\sigma_{(0,1)}=\sigma_{z}$. The evaluations of the expectation values and nonlocal violations for the qudit Bell operators in Eq. (14) and Eq. (20) do not straightforwardly generalize to qubits using the above Wigner function. The reason is that the assignments of an lhv model are different from qudit Pauli operators to qubit Pauli operators if one measures a complete set of operators generating mutually unbiased bases. However, in case where $U_{\nu}=T$, the equivalent Bell operator for qubits in Eq. (14) is
$\mathcal{B}^{\phantom{\prime}}_{2}=\mathds{1}\otimes\mathds{1}+\sigma_{z}\otimes\sigma_{z}-\frac{1}{\sqrt{2}}\Big{(}T\sigma_{y}T^{\dagger}\otimes\sigma_{y}+T\sigma_{x}T^{\dagger}\otimes\sigma_{y}+T\sigma_{y}T^{\dagger}\otimes\sigma_{x}-T\sigma_{x}T^{\dagger}\otimes\sigma_{x}\Big{)}\,,$ (34)
@@ Appendix C Bounds on Gross’ Wigner function
$\mathrm{Tr}\left(\rho\mathcal{B}\right)$ $=d^{2}\sum_{u_{1},u_{2}\in\mathbb{Z}_{d}^{2}}W^{\nu}_{u_{1},u_{2}}W_{u_{1},u_{2}}\left((U_{\nu}\otimes\mathds{1})^{\dagger}\rho\,(U_{\nu}\otimes\mathds{1})\right)$
$\leq d^{2}\max_{u_{1},u_{2}\in\mathbb{Z}_{d}^{2}}|W^{\nu}_{u_{1},u_{2}}|\sum_{u_{1},u_{2}\in\mathbb{Z}_{d}^{2}}|W_{u_{1},u_{2}}\left((U_{\nu}\otimes\mathds{1})^{\dagger}\rho\,(U_{\nu}\otimes\mathds{1})\right)|\,.$ (37)